### v1

# Self-intersection Number of Negative Curves on Fermat Surfaces

Abstract.
We give an explicit formula for the self-intersection number of negative curves on Fermat surfaces. The formula gives us hints to either prove or disprove the Bounded Negativity Conjecture for the Fermat surfaces.
Key words and phrases:
Bounded Negativity Conjecture, Fermat surface, intersection multiplicity, Puiseux series, resultant
2010 Mathematics Subject Classification:
Primary 14C17, Secondary 14H20.
1. Introduction
Let $\mathbb{P}^{3}$ be the three-dimensional complex projective space, with homogeneous coordinate ring $S=\mathbb{C}[x,y,z,w]$. The Fermat surface of degree $d$, denoted as $X_{d}$, is the smooth hypersurface in $\mathbb{P}^{3}$ defined by the equation:
$x^{d}+y^{d}+z^{d}+w^{d}=0.$
The main objective in this paper is to establish the following theorem.
Theorem 1.1 (Theorem 4.5).
Let $X_{d}\subset\mathbb{P}^{3}$ be the Fermat surface defined by $x^{d}+y^{d}+z^{d}+w^{d}=0$ for $d\geq 5$, and $C$ a reduced and irreducible curve on $X_{d}$ with $C^{2}<2-d<0$. Then
(1) There exists an irreducible homogeneous polynomial $g(x,y,z)$ which vanishes on $C$.
Denote by $e=\deg g$. Define two curves $E$ and $D$ in $\mathbb{P}^{2}$ by
$E:\ g=0\quad\mathrm{and}\quad D:x^{d}+y^{d}+z^{d}=0.$
(2) There exists $k\geq 2$ and $k|d$ such that
$C^{2}$ $=$ $\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)$
$+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E),$
where $\mu_{p}(E)$ denotes the Milnor number of $E$ at $p$.
In particular, when $d$ is prime, then $k=d$ and
$C^{2}=e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C).$
This result is closely connected with an old folklore conjecture among experts in algebraic surface theory, commonly known as the Bounded Negativity Conjecture. We formulate this conjecture as follows.
Conjecture 1.2 (Bounded Negativity Conjecture).
For any given smooth complex projective surface $X$, there exists a number $B_{X}\geq 0$ which depends only on $X$ such that for any reduced curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $-B_{X}$, i.e., $C^{2}\geq-B_{X}$.
The origins of this conjecture are unclear, but it can at least be traced back to the Italian mathematician Federigo Enriques (1871-1946), and it has been an open problem for one hundred years. For a more comprehensive overview , we refer to [3, p.1878].
An alternative formulation of the Bounded Negativity Conjecture is as follows.
Conjecture 1.3 (Bounded Negativity Conjecture for integral curves).
For any given smooth complex projective surface $X$, there exists a number $b_{X}\geq 0$ which depends only on $X$ such that for any reduced and irreducible curve $C$ on $X$, the self-intersection number $C^{2}$ is bounded below by $-b_{X}$, i.e., $C^{2}\geq-b_{X}$.
Indeed, as proven in [3, Prop.5.1, p.1891], we can take the number $B_{X}$ in Conjecture 1.2 as $B_{X}=\left(\rho(X)-1\right)b_{X}$, where $\rho(X)$ is the Picard number of $X$. We will refer to $b_{X}$ as the negativity bound for $X$.
Our result in Theorem 1.1 implies that under some conditions, the Bounded Negativity Conjecture is true for the Fermat surface $X_{d}$.
Corollary 1.4 (Corollary 4.6).
The Bounded Negativity Conjecture holds on the Fermat surface $X_{d}$ if one of the following conditions is satisfied:
$\liminf_{e\to\infty}\inf_{E}\left(e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)\right)\geq-\lambda$
for some constant $\lambda>0$ which may depend on $d$, where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
$\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1,$
where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
On the other hand, the formula in Theorem 1.1 also gives us hint to construct counterexamples for the Bounded Negativity Conjecture on Fermat surfaces; see Section 5.
In the sequel, we will only investigate the self-intersection number of reduced and irreducible curves. For convenience, we denote $C$ such a curve on a given smooth projective surface.
It is evident that the Bounded Negativity Conjecture holds for $X$ with $-K_{X}\geq 0$, as can be seen by applying the adjunction formula:
$K_{X}\cdot C+C^{2}=2p_{a}(C)-2\Longrightarrow C^{2}\geq-2.$
In particular, for $d\leq 4$, the bounded negativity of $X_{d}$ follows from this argument, since the canonical divisor of $X_{d}$ is $K_{X_{d}}=(d-4)H$ where $H$ is the hyperplane section. However, for $d\geq 5$, the canonical divisor $K_{X_{d}}$ is ample and $X_{d}$ is of general type. To the author’s knowledge, there are currently no effective methods available to deal with the Bounded Negativity of such surfaces.
Another natural approach to establishing the Bounded Negativity for a surface is to show that there are only finitely many negative curves. However, it is proved in [3, Thm.4.1] that for any $m>0$, there exists a smooth projective surface which admits infinitely many smooth irreducible curves of self-intersection number $-m$. For the Fermat surface $X_{d}$, we do not know whether the number of negative curves on $X_{d}$ is finite or not.
A third potential method to prove the Bounded Negativity Conjecture is to establish the boundedness of the geometric genera of reduced and irreducible curves. In fact, the weak Bounded Negativity Conjecture, proven in [10], states that on a given smooth projective surface $X$, the inequality
$C^{2}\geq-b_{X,g(C)}$
holds for any reduced and irreducible curve $C$, where $b_{X,g(C)}$ is a constant depending only on $X$ and the geometric genus $g(C)$ of $C$. Consequently, the Bounded Negativity Conjecture would follow if we could establish a bound for the genera $g(C)$. Nevertheless, such a bound does not exist for almost all surfaces. For the Fermat surface $X_{d}$, consider a generic curve $C$ in the linear series $|kH|$, then $C$ is a smooth irreducible curve on $X_{d}$ with geometric genus
$g(C)=\frac{1}{2}(K_{X_{d}}\cdot C+C^{2})+1=\frac{1}{2}\left(d(d-4)k+dk^{2}\right)+1,$
thus the genus $g(C)$ goes to infinity as $k$ tends to infinity. Therefore, the Bounded Negativity Conjecture for $X_{d}$ cannot be proved by bounding the genera of reduced and irreducible curves.
Other strategies to prove the Bounded Negativity Conjecture include investigating the non-isomorphic endomorphisms or establishing dominance by other surfaces with bounded negativity [3, Prop.2.1], exploring the birational invariance of bounded negativity and the $H$-constant [2], using the equivalence between bounded negativity and the boundedness of the denominators in the Zariski decomposition [4], discussing the weighted Bounded Negativity Conjecture [11, 8], and studying the connections between the Bounded Negativity Conjecture and other unsolved problems, such as SHGH conjecture or the bounded cohomology conjecture [1, 7]. However, despite providing invaluable insights in understanding the Bounded Negativity Conjecture, these methods either deal with only very special cases, e.g., rational surfaces or K3 surfaces, or just offer general frameworks that reduce the conjecture to other open questions. They rely on general formulas, such as the adjunction formula and Riemann-Roch formula, to establish bounds for the intersection number $C^{2}$ of a reduced and irreducible curve $C$, without using specific information about $C$, such as its equation.
Can we estimate $C^{2}$ from the equation of the curve $C$? Can we give a concrete example of a surface with the Bounded Negativity by doing explicit computation through the defining equations of the surface?
To explore these questions, we consider smooth surfaces in $\mathbb{P}^{3}$, and give the following conjecture.
Conjecture 1.5.
For any $d\geq 1$, there exists a constant $c_{d}$ that depends only on $d$ such that all smooth surfaces of degree $d$ in $\mathbb{P}^{3}$ has negativity bound $c_{d}$.
Since all the degree $d$ surfaces in $\mathbb{P}^{3}$ are contained in a connected family, they can be deformed into one another. We further conjecture that the bounded negativity property is invariant under deformations.
Conjecture 1.6.
Suppose $X$ and $Y$ are two smooth complex projective surfaces that are in a single family of complex projective surfaces, then the Bounded Negativity Conjecture is true for $X$ if and only if it is true for $Y$.
Hence, to establish the Bounded Negativity Conjecture for surfaces in $\mathbb{P}^{3}$, it suffices to prove its validity for a specific surface. Our Theorem 1.1 is proved for this purpose.
The formulas in Theorem 1.1 show how the Bounded Negativity Conjecture is connected with the singularity theory of curves. The key term in the expression of the self-intersection number is the sum $\sum\limits_{p\in D\cap E}\mu_{p}(E)$. It is notable that this sum is taking over the intersection points of $D$ and $E$, not all the singular points of $E$.
To conclude the introduction, we give an outline of the proof of Theorem 1.1. Given a negative curve $C$ on $X_{d}$. First, we represent the Fermat surface $X_{d}$ as a branched Galois cover of $\mathbb{P}^{2}$:
$\rho:\ X_{d}\to\mathbb{P}^{2}\quad[x,y,z,w]\mapsto[x,y,z]$
Let $E=\rho(C)$ and $\mathscr{C}=\rho^{*}E$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$; denote by $g$ its defining equation. It is clear that $g$ can be seen as a homogeneous polynomial on $\mathbb{P}^{3}$ which vanishes on $C$. Second, we prove that $\mathscr{C}$ is a reduced curve on $X_{d}$; see Proposition 4.4. Moreover, $\mathscr{C}$ is of the following form
$\mathscr{C}=C_{1}+\cdots+C_{k},$
with $C_{i}^{2}=C$ for all $i=1,\cdots,k$. This result holds because $C_{i}$ lie in the orbit of $C$ under the Galois group action for $\rho$. It follows that
$kC^{2}=d(\deg g)^{2}-\sum_{i\neq j}C_{i}\cdot C_{j}.$
Hence, to compute $C^{2}$, we need to calculate the sum on the right hand side. Third, the intersection number can be calculated in a local way at any intersection point. Locally, the curve $\mathscr{C}$ can be represented as the curve in $(\mathbb{C}^{3},0)$ defined by
$h(u,v)=0,\ l(u,v)+\omega^{d}=0.$
We will first solve $v$ as Puiseux series of $u$ from the first equation, and then use $l\left(u,v(u)\right)+\omega^{d}=0$ to investigate the intersection multiplicities between different irreducible components of $\mathscr{C}$. This is the most technical part in this paper and is done in Proposition 4.4 and Section 6.
The author would like to thank the organizers of International Symposium on Singularities and Applications, Dec. 9-13, 2024 in Sanya, China. The stimulating atmosphere of the conference inspired him to conceive the idea of performing computation using Puiseux series. The author was partially supported by the grant YSBR-001 and is supported by NSFC No. 12301100.
2. Curves on surfaces of Sebastiani-Thom type
Assume that $X\subset\mathbb{P}^{3}$ be a smooth surface defined by the following equation:
$X:f(x,y,z)+w^{d}=0,$
where $f$ is a binary form of degree $d$. Such a surface is said to be of Sebastiani-Thom type. The Fermat surface is a typical example. Let $C$ be a reduced and irreducible curve on $X$.
In this section, we show that $C$ is “almost” defined by a homogeneous polynomial on $X$ in the following result.
Proposition 2.1.
Let $X:f(x,y,z)+w^{d}=0$ be a smooth surface in $\mathbb{P}^{3}$ and $C$ an reduced and irreducible curve on $X$. Then there exists an irreducible form $g(x,y,z)$ satisfying the following properties:
  1. (i)
    The divisor on $X$ cut out by $g$ can be written as
    $(g)=m(C_{1}+\cdots+C_{k}),$
    where $C_{1}=C,C_{2},\cdots,C_{k}$ are distinct reduced and irreducible curves and $m\geq 1$ is an integer.
  2. (ii)
    The self-intersection numbers $C_{i}^{2},i=1,\cdots,k$ are all equal to $C^{2}$.
Proof.
Consider the following projection morphism
$\rho:X\to\mathbb{P}^{2},\ [x,y,z,w]\mapsto[x,y,z].$
It is a branched Galois covering map with Galois group $\mathbb{Z}_{d}=\mathbb{Z}/d\mathbb{Z}$, and the Galois action is given by
$a\cdot[x,y,z,w]=[x,y,z,e^{2\pi\sqrt{-1}a/d}w],\ \forall a\in\mathbb{Z}_{d}.$
For the curve $C\subset X$, we consider its image under the morphism $\rho$, denoted by $E=\rho(C)$. Then $E$ is a reduced and irreducible curve in $\mathbb{P}^{2}$.
Let $g(x,y,z)$ be an irreducible defining equation for $E$. Then the inverse image $\rho^{-1}(E)$ is exactly the curve cut out by $g$ on $X$. We may see $g$ as a homogeneous polynomial on $X$ and write the divisor $(g)$ cut out by $g$ on $X$ as
$(g)=m_{1}C_{1}+\cdots+m_{k}C_{k},\ C_{1}=C,$
where $C_{1},\cdots,C_{k}$ are reduced and irreducible curves on $X$ and $m_{1},\cdots,m_{k}\geq 1$.
Since the action $\mathbb{Z}_{d}$ for the morphism $\rho$ is a Galois action, we see that $\rho^{-1}(E)$ is exactly the union of the curves in the $\mathbb{Z}_{d}$-orbit of $C$. It follows that $C_{i},i>1$ are all isomorphism to $C$ by an automorphism of $X$. Hence $C_{i}^{2}=C^{2}$ for all $i=1,\cdots,k$; and (i) follows.
Moreover, we may see $g$ as a $\mathbb{Z}_{d}$-invariant polynomial. For each pair $i\neq j$, there exists an element $\alpha$ in the Galois group $\mathbb{Z}_{d}$ such that $\alpha\cdot C_{i}=C_{j}$, and then
$m_{i}=\mathrm{ord}_{C_{i}}(g)=\mathrm{ord}_{\alpha\cdot C_{i}}(\alpha\cdot g)=\mathrm{ord}_{C_{i}}(g)=m_{j}.$
Hence (ii) is true. ∎
3. Strategy of Proof
Let $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$ is the Fermat surface in $\mathbb{P}^{3}$ and $C\subset X_{d}$ is a reduced and irreducible curve. We shall investigate the lower bound of the self-intersection number $C^{2}$.
By Proposition 2.1, there exists an irreducible polynomial $g(x,y,z)$ which cuts out a divisor on $X$ of the following form
$(g)=m(C_{1}+\cdots+C_{k}),\ C_{1}=C$
and $C_{i}^{2}=C^{2}$ for $i=1,\cdots,k$.
To investigate the Bounded Negativity Conjecture for $X_{d}$, we can safely assume that $k>1$; otherwise, $m^{2}C^{2}=d(\deg g)^{2}$ and thus $C^{2}\geq 0$. Therefore, we assume $k\geq 2$ in the sequel.
For convenience, denote by $D$ the curve defined by $x^{d}+y^{d}+z^{d}=0$ and $E$ the curve defined by $g=0$ in $\mathbb{P}^{2}$. Namely,
$D:x^{d}+y^{d}+z^{d}=0\qquad\mathrm{and}\qquad E:g(x,y,z)=0.$
Denote $e=\deg g=\deg E$.
Denote $\mathscr{C}$ the scheme on $X_{d}$ defined by $g=0$:
$\mathscr{C}:g(x,y,z)=0.$
Then $C_{1},\cdots,C_{k}$ are the irreducible components of $\mathscr{C}$.
3.1. Strategy
From the equation
$(g)=m(C_{1}+\cdots+C_{k}),$
we get
$de^{2}$ $=$ $m^{2}(C_{1}+\cdots+C_{k})^{2}$
$=$ $m^{2}\sum_{i=1}^{k}C_{i}^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j}$
$=$ $m^{2}kC^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j},$
where $C_{i}\cdot C_{j}$ is the intersection number on $X_{d}$.
To investigate the Bounded Negativity Conjecture on $X_{d}$, we need to obtain a lower bound for $C^{2}$. From the formula above, we are led to give an upper bound for the sum
$\sum_{i\neq j}C_{i}\cdot C_{j}.$
The result will be a function in $e$ where $d$ is viewed as a constant, and so is the difference
$de^{2}-m^{2}\sum_{i\neq j}C_{i}\cdot C_{j}.$
If this function is bounded below for $e\geq 1$, then the Bounded Negativity Conjecture for $X_{d}$ will follow.
The main part of this paper is to calculate the intersection numbers by using local computation.
For an arbitrary intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of some $C_{i}$ and $C_{j}$ with $i\neq j$, we will give an explicit formula for the sum
$\sum_{i\neq j}I_{P}(C_{i},C_{j}),$
where $I_{P}(C_{i},C_{j})$ is the intersection multiplicity (that will be explored in detail in Section 6).
Since we consider all curves $C$ on $X_{d}$, we see $g$ as an abstract polynomial. From $P\in X_{d}$, we have that $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}+w_{0}^{d}=0$, hence not all $x_{0},y_{0},z_{0}$ are zero. By the symmetry of the coordinates $x,y,z$ in the equation $X_{d}:x^{d}+y^{d}+z^{d}+w^{d}=0$, we may without loss of generality assume that $x_{0}\neq 0$. To decide where the intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ may situate on $X_{d}$, we propose the following Lemma which will be proved later.
Lemma 3.2.
Following the notations above, for any point $Q$ lying on at least two irreducible components of $\mathscr{C}$, we have
$w(Q)=0,$
where $w$ is the fourth homogeneous coordinate for $\mathbb{P}^{3}$. Moreover, $Q\in C_{i}$ for all $i=1,\cdots,k$, i.e., $Q$ lies on every irreducible component of $\mathscr{C}$.
In particular, for the point $P=[x_{0},y_{0},z_{0},w_{0}]$ above, we have $w_{0}=0$ and $P\in C_{i}$ for all $i\leq k$.
Let $p=\rho(P)=[x_{0},y_{0},z_{0}]$. Then $p\in E\cap D$.
Recall that $w_{0}=0$ and we assumed $x_{0}\neq 0$. To calculate the intersection multiplicities at $P$, we introduce affine coordinates
$u=\frac{y}{x}-\frac{y_{0}}{x_{0}},\qquad v=\frac{z}{x}-\frac{z_{0}}{x_{0}},\qquad\omega=\frac{w}{x}.$
Then the curve $\mathscr{C}:g=0$ on $X_{d}$ is defined in $\mathbb{C}^{3}=\{x\neq 0\}\subset\mathbb{P}^{3}$ by
$h(u,v)=0,\quad l(u,v)+\omega^{d}=0,$
where
$h(u,v)=g\left(1,u+\frac{y_{0}}{x_{0}},v+\frac{z_{0}}{x_{0}}\right)$
and
$l(u,v)=1+\left(u+\frac{y_{0}}{x_{0}}\right)^{d}+\left(v+\frac{z_{0}}{x_{0}}\right)^{d}.$
Moreover, since $x_{0}^{d}+y_{0}^{d}+z_{0}^{d}=0$ and $x_{0}\neq 0$, at least one of $y_{0},z_{0}$ does not vanish. We may without loss of generality assume that $z_{0}\neq 0$. Then
$\frac{\partial\left(l(u,v)+\omega^{d}\right)}{\partial v}(0,0,0)=d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}\neq 0.$
Therefore, we can locally express $v$ as an analytic function of $(u,\omega)$ by applying the implicit function theorem to the equation $l(u,v)+\omega^{d}=0$. Then the local equation for $\mathscr{C}$ is $h\left(u,v(u,\omega)\right)=0$. By analyzing this local equation, we can investigate the local branches and compute the local intersection multiplicities. This is the usual way to analyze the curve $\mathscr{C}$.
Our method is, however, first solving $v$ from the equation $h(u,v)=0$ by using the Puiseux series method (that will be recalled in Section 6), say $v=s(u)$. Then using the equation $l(u,s(u))+\omega^{d}=0$ to discuss the properties of the curve $\mathscr{C}$ and calculate the intersection multiplicities. To this end, we need that $h$ is $v$-general (the related notions will be discussed in Section 6).
At this point, we specially consider the exceptional case where $h$ is not $v$-general, or equivalently, $h(0,v)\equiv 0$; the $v$-general case will be investigated in subsequent sections.
Recall that $h$ is transformed from an irreducible polynomial $g$, thus $h$ is an irreducible polynomial in $u,v$. If $h$ is not $v$-general, then $u$ is a factor of $h$ and so $h$ is a constant multiple of $u$ by irreducibility. Back to $g$, we see that $g$ is a linear function $y-ax$ (up to multiplicative factor). Therefore, the curve $\mathscr{C}$ is defined as
$\mathscr{C}:y=ax,\ (1+a^{d})x^{d}+z^{d}+w^{d}=0.$
If $a^{d}+1\neq 0$, then $\mathscr{C}$ is itself irreducible and hence it must be equal to $C$. Thus, $C^{2}=d$ in this case.
If $a^{d}+1=0$, then $\mathscr{C}$ splits into $d$ lines. Let
$L_{i}:y=ax,\ z=e^{(1+2i)\pi\sqrt{-1}/d}w,$
for $i=1,\cdots,d$. Then
$(g)=L_{1}+\cdots+L_{d}.$
It is straightforward that $L_{i}\cdot L_{j}=1$ for $i\neq j$. Hence, by the above discussions, we have
$d=dC^{2}+d(d-1),$
and so $C^{2}=2-d$.
3.3. Proof of Lemma 3.2
Recall that we used the degree $d$ branched Galois cover $\rho:X_{d}\to\mathbb{P}^{2}$ defined by $[x,y,z,w]\mapsto[x,y,z]$. The ramification locus of $\rho$ is the closed subset of $X_{d}$ defined by the equation $w=0$.
If $Q$ is a point on at least two components of $\mathscr{C}$, then #$\#\left(\rho^{-1}(\rho(Q))\right)<d$, where #$\#A$ means the number of points of a set $A$. It follows that $Q$ lies on the ramification locus of $\rho$. Thus $w(Q)=0$ as desired.
Furthermore, denote $q=\rho(Q)$, then $\rho^{-1}(q)=\{Q\}$. For all $i\leq k$, the restriction morphism $\rho|_{C_{i}}:C_{i}\to E$ is a finite morphism and thus $\rho^{-1}(q)\cap C_{i}\neq\emptyset$. But $\rho^{-1}(q)$ consists of only one point $Q$, it follows that $Q\in C_{i}$. We are done.
4. Self-intersection number for curves on Fermat surfaces
Following the notations above, we will give an explicit formula for the self-intersection number $C^{2}$.
4.1. Position of intersection points
From Lemma 3.2, we see that any intersection point $Q$ of two irreducible components of $\mathscr{C}$ satisfies $w(Q)=0$. It follows that $q=\rho(Q)$ is an intersection point of $D$ and $E$. The converse is also true.
Lemma 4.2.
If $q=[x,y,z]$ lies on both $D$ and $E$, then $Q=[x,y,z,0]$ lies on all of $C_{i},i\leq k$.
Hence, $Q\in C_{i}\cap C_{j}$ for $i\neq j$ if and only if $q\in D\cap E$.
Proof.
By definition, $\rho^{-1}(q)=Q$ and $Q\in\mathscr{C}$. Since each $C_{i}$ is mapped onto $E$, the intersection $\rho^{-1}(q)\cap C_{i}$ is nonempty. Hence $Q\in C_{i}$ as desired.
The last assertion follows immediately by Lemma 3.2. ∎
4.3. Intersection multiplicities
Recall that we have the following equation
$de^{2}=m^{2}kC^{2}+m^{2}\sum_{i\neq j}C_{i}\cdot C_{j}.$
To give a lower bound for $C^{2}$, we need to estimate the sum on the right hand side. Moreover, by definition, we have
$\sum_{i\neq j}C_{i}\cdot C_{j}=\sum_{\stackrel{{\scriptstyle i\neq j}}{{Q\in C_{i}\cap C_{j}}}}I_{Q}(C_{i},C_{j}).$
Hence, it suffices to do computation for the sum
$\sum_{i\neq j}I_{P}(C_{i},C_{j})$
where $P=[x_{0},y_{0},z_{0},0]$ is chosen above. Our main technical results are as follows.
Proposition 4.4.
With the notations above, we have
  1. (A)
    $m=1$; in other words, $\mathscr{C}$ is a reduced scheme.
  2. (B)
    The following equation holds:
    $\mu_{P}(\mathscr{C})=(d-1)i_{p}(D,E)+d\mu_{p}(E)-d+1.$
  3. (C)
    We have
    $\sum_{i\neq j}I_{P}(C_{i},C_{j})=(d-1)i_{p}(D,E)+d\mu_{p}(E)-k\mu_{P}(C)-(d-k).$
    and thus
    $\sum_{i\neq j}C_{i}\cdot C_{j}=d(d-1)e+d\sum_{p\in D\cap E}\mu_{p}(E)-k\sum_{p\in D\cap E}\mu_{\rho^{-1}(p)}(C)-(d-k)\#(D\cap E).$
The proof will be presented in Section 6.
To investigate the Bounded Negativity Conjecture for Fermat surfaces, we only need to consider negative curves $C$. Then the curve $\mathscr{C}$ necessarily have $k\geq 2$ irreducible components since otherwise $C^{2}$ would be nonnegative. Moreover, $k$ divides $d$ since the Galois action for $\rho$ is a cyclic action of order $d$. To ensure that the polynomial $h$ is $v$-general, we only need to assume $C^{2}\neq 2-d$. Therefore, combining all of our discussions above, we get the following theorem.
Theorem 4.5.
Let $X_{d}\subset\mathbb{P}^{3}$ be the Fermat surface defined by $x^{d}+y^{d}+z^{d}+w^{d}=0$ for $d\geq 5$, and $C$ a reduced and irreducible curve on $X_{d}$ with $C^{2}<2-d<0$. Then
(1) There exists an irreducible homogeneous polynomial $g(x,y,z)$ which vanishes on $C$.
Denote by $e=\deg g$. Define two curves $E$ and $D$ in $\mathbb{P}^{2}$ by
$E:\ g=0\quad\mathrm{and}\quad D:x^{d}+y^{d}+z^{d}=0.$
(2) There exists $k\geq 2$ and $k|d$ such that
$C^{2}$ $=$ $\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)$
$+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E).$
In particular, when $d$ is prime, then $k=d$ and
$C^{2}=e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C).$
As a corollary, we get two sufficient conditions for the Bounded Negativity Conjecture to hold on Fermat surfaces.
Corollary 4.6.
With the notations above, the Bounded Negativity Conjecture holds on the Fermat surface $X_{d}$ if one of the following conditions is satisfied:
$\liminf_{e\to\infty}\inf_{E}\left(e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)\right)\geq-\lambda$
for some constant $\lambda>0$ which may depend on $d$, where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
$\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}<1,$
where $E$ ranges over all irreducible projective plane curves of degree $e$ and $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
Proof.
By our computations before, it is clear that (a) implies the Bounded Negativity Conjecture. Indeed, if (a) holds, there exists $e_{0}>d+1$ such that
$\inf_{E}\left(e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)\right)\geq-(\lambda+1)$
for $e\geq e_{0}$. On the other hand, the geometric genus $g(E)$ of $E$ is given by
$g(E)=\frac{(e-1)(e-2)}{2}-\sum_{p\in E}\left(\mu_{p}(E)+r_{p}-1\right)$
by [12, Cor.7.1.3], where $r_{p}$ is the number of local branches of $E$ at $p$. If follows that
$\sum_{p\in D\cap E}\mu_{p}(E)\leq\sum_{p\in E}\mu_{p}(E)\leq(e-1)(e-2)$
Hence, for $e<e_{0}$, we have
$e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)$$\geq$ $e^{2}-(d-1)e-(e-1)(e-2)$
$=$ $(4-d)e-2$
$\geq$ $-(d-4)e_{0}-2.$
Therefore, we may take
$b_{X_{d}}=\max\{\lambda+1,(d-4)e_{0}+2).$
The condition (b) implies (a), as can be seen as follows: if (b) holds, then for $e$ sufficiently large, we have
$e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)\geq e^{2}-(d-1)e-(1-\alpha)e^{2}=\alpha e^{2}-(d-1)e.$
for some $\alpha>0$. The result follows by noting that
$\alpha e^{2}-(d-1)e=\alpha\left(e-\frac{d-1}{2\alpha}\right)^{2}-\frac{(d-1)^{2}}{4\alpha}\geq-\frac{(d-1)^{2}}{4\alpha}.$
5. Discussions and Problems
In this section, we investigate some cases in which we are able to estimate the intersection number $C^{2}$ which is equal to
$\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E).$
by Theorem 4.5.
The first key obstacle for us to get the Bounded Negativity comes from the term
$\sum_{p\in D\cap E}\mu_{p}(E)$
which may have order $e^{2}$ for $e$ sufficiently large. Moreover, this problem arises from two possible causes: (i) $D\cap E$ contains many singular points of $E$; (ii) $\mu_{p}(E)$ is large for $p\in D\cap E$.
5.1. The case #$\#(D\cap E)=de$.
We first consider the case where $D\cap E$ contains as many points as possible. In this case, #$\#(D\cap E)=de$. Then for each $p\in E$, we have $\mu_{p}(E)=0$ by [6, Cor.2.6.8]. Therefore,
$\sum_{p\in D\cap E}\mu_{p}(E)=0.$
5.2. The case $\mu_{p}(E)$ is large for some $p\in D\cap E$.
In the extreme case, #$\#(D\cap E)=1$ and $D\cap E$ is the only singular point of $E$; moreover, $E$ is a rational cuspidal curve and $C$ is smooth. Then
$\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C)+\left(\frac{d}{k}-1\right)\#(D\cap E)$
$=$ $\frac{d}{k}e^{2}-\frac{d(d-1)}{k}e-\frac{d}{k}(e-1)(e-2)+\left(\frac{d}{k}-1\right)$
$=$ $\frac{d}{k}\left((4-d)e-1\right)-1,$
which can be very negative for $e$ sufficiently large.
Hence we propose the following problem:
Problem 5.3.
Does there exist a reduced and irreducible homogeneous polynomial $g_{e_{i}}$ for a sequence $e_{i}\to\infty$ satisfying the following conditions?
  1. (i)
    The projective plane curve $V(g_{e_{i}})$ is a rational cuspidal curve with only one singular point which lies on the curve $x^{d}+y^{d}+z^{d}=0$; and
  2. (ii)
    The curve $\mathscr{C}_{e_{i}}$ in $\mathbb{P}^{3}$ defined by $g_{e_{i}}(x,y,z)=0,x^{d}+y^{d}+z^{d}+w^{d}=0$ is a reducible curve with smooth irreducible components.
The the answer is positive, then taking an irreducible component of $\mathscr{C}_{e_{i}}$ would disprove the Bounded Negativity Conjecture for $X_{d}$.
5.4. The case $C$ is a smooth rational curve.
If $C$ is a rational curve, then necessarily $k=d$ no matter whether $d$ is prime or not, since in this case, the finite morphism $\rho:C\to D$ reduces to a birational morphism. Therefore, we have
$C^{2}=e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E)+\sum_{[x,y,z]\in D\cap E}\mu_{[x,y,z,0]}(C).$
To remove the influence caused by the terms $\mu_{[x,y,z,0]}(C)$, we assume in addition that $C$ is smooth. Then
$C^{2}=e^{2}-(d-1)e-\sum_{p\in D\cap E}\mu_{p}(E).$
Hence we propose the following problem:
Problem 5.5.
Does there exist a reduced and irreducible homogeneous polynomial $g_{e_{i}}$ for a sequence $e_{i}\to\infty$ satisfying the following conditions?
  1. (i)
    The projective plane curve $V(g_{e_{i}})$ is a rational curve; and
  2. (ii)
    The curve $\mathscr{C}_{e_{i}}$ in $\mathbb{P}^{3}$ defined by $g_{e_{i}}(x,y,z)=0,x^{d}+y^{d}+z^{d}+w^{d}=0$ is a reducible curve with smooth rational irreducible components; and
  3. (iii)
    The sequence
    $e_{i}^{2}-(d-1)e_{i}-\sum_{p\in D\cap V(g_{e_{i}})}\mu_{p}(V(g_{e_{i}}))$
    is not bounded from below, where $D$ is the projective plane curve defined by $x^{d}+y^{d}+z^{d}=0$.
The the answer is positive, then the Bounded Negativity Conjecture will not hold for $X_{d}$.
5.6. An invariant
From Corollary 4.6 (b), we are led to consider the following invariant
$\limsup_{e\to\infty}\sup_{E}\frac{\sum\limits_{p\in D\cap E}\mu_{p}(E)}{e^{2}}.$
More generally, given any reduced homogeneous polynomial $f$ of degree $d$, we define
$I_{f}=\limsup_{e\to\infty}\sup_{g\ \mathrm{irreducible}}\frac{\sum\limits_{p\in V(f)\cap V(g)}\mu_{p}\left(V(g)\right)}{e^{2}},$
where $V(g)$ is the projective plane curve defined by $g$.
Problem 5.7.
  1. (1)
    If $f=x^{d}+y^{d}+z^{d}$, is it true that $I_{f}<1$?
  2. (2)
    More generally, if $f$ defines a smooth projective surface in $\mathbb{P}^{2}$, does it hold that $I_{f}<1$?
  3. (3)
    Does the invariant $I_{f}$ depends on the equation of $f$, or at least the degree $d$ of $f$?
6. Puiseux series, resultants, and intersection multiplicities
In this section, we first recall basic notions concerning Puiseux series and resultants, and then use them to prove properties of intersection multiplicities of plane curve germs. Finally, we give the proof of Proposition 4.4.
6.1. Puiseux series
We follow the notations in [6, Sec.1.2, p.17]. The ring $\mathbb{C}[[u]]$ denotes the set of formal power series, and $\mathbb{C}\{u\}$ the analytic local ring of $(\mathbb{C},0)$. Denote by $\mathbb{C}((u))$ the field of fractions of $\mathbb{C}[[u]]$. A typical element of $\mathbb{C}((u))$ can be represented as
$\sum_{i\geq r}a_{i}u^{i},$
for some $r\in\mathbb{Z}$ and $a_{i}\in\mathbb{C}$.
The field of fractionary power series, denoted by $\mathbb{C}\langle\langle u\rangle\rangle$ is the direct limit of the fields $\mathbb{C}((u^{\frac{1}{n}}))$, where the morphisms in the directed system are
$\mathbb{C}((u^{\frac{1}{n}}))\to\mathbb{C}((u^{\frac{1}{n^{\prime}}})),\quad n|n^{\prime}\mathrm{\ and\ }\sum a_{i}u^{\frac{i}{n}}\mapsto\sum a_{i}x^{\frac{in^{\prime}/n}{n^{\prime}}}.$
Let
$s=\sum_{i\geq r}a_{i}u^{\frac{i}{n}}$
be a fractionary power series, we define the order of $s$ to be
$o_{u}(s)=\begin{cases}\infty,&\mathrm{if\ }s=0;\\ \frac{\min\{i:a_{i}\neq 0\}}{n},&\mathrm{if\ }s\neq 0.\end{cases}$
Fractionary power series $s$ with $o_{u}(s)>0$ will be called Puiseux series. After a simultaneous reduction of all the fractional exponents effectively appearing in $s$, we may assume that $n$ and $\gcd\{i:a_{i}\neq 0\}$ are comprime, and then we call $n$ the polydromy order of $s$, denoted by $\nu(s)$. Moreover, we say that $s$ is a convergent fractionary power series if the formal power series $s(t^{n})$ has a positive convergence radius.
Suppose $f\in\mathbb{C}[[u,v]]$ is a formal power series satisfying $f(0,0)=0$. We say a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a $v$-root (or simply, root) of $f$ is $f(u,s(u))=0$. Furthermore, we say that $f$ is $v$-general of order $k$, if
$f(0,v)=c_{k}v^{k}+c_{k+1}v^{k+1}+\cdots,$
for some constant $c_{k}\neq 0$.
For convenience, for $s,s^{\prime}\in\mathbb{C}\langle\langle u\rangle\rangle$, we use $s\simeq s^{\prime}$ to denote the relation
$s=(a+s^{\prime\prime})s^{\prime}$
for some $a\in\mathbb{C}^{*}$ and $s^{\prime\prime}$ a Puiseux series. Note that if $s\simeq s^{\prime}$, then $o_{u}(s)=o_{u}(s^{\prime})$.
Some basic results that we will use in the theory of Puiseux series are summarized as follows; we refer to [6, Chap.1] for the proof.
Proposition 6.2.
The following statements are true:
  1. (i)
    The field $\mathbb{C}\langle\langle u\rangle\rangle$ is algebraically closed.
  2. (ii)
    If $f\in\mathbb{C}\{u,v\}$ is a convergent power series and a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a root of $f$, then $s$ is convergent.
  3. (iii)
    If $f\in\mathbb{C}\{u,v\}$ is $v$-general of order $b$ and a Puiseux series $s\in\mathbb{C}\langle\langle u\rangle\rangle$ is a root of $f$ such that $\nu(s)=b$, then
    $\prod_{\alpha=1}^{b}\left(v-s(e^{2\pi\sqrt{-1}\alpha/b}u)\right)\mathrm{\ divides\ }f\mathrm{\ in\ }\mathbb{C}\{u,v\}.$
    Moreover, if, in addition, $f$ is irreducible, then
    $f=c(u,v)\prod_{\alpha=1}^{b}\left(v-s(e^{2\pi\sqrt{-1}\alpha/b}u)\right)$
    for an invertible element $c(u,v)$ in $\mathbb{C}\{u,v\}$.
6.3. Resultants
We recall the definition and basic properties of resultants; all the results are well known. The interested reader may consult [5, Sec.1.4].
Let $K$ be a field of characteristic zero, which will be taken as $\mathbb{C}\langle\langle u\rangle\rangle$ in our discussion about intersection multiplicities. Let $f,g$ are two polynomials in $K[v]$ of the following form:
$f=a_{0}v^{d}+a_{1}v^{d-1}+\cdots+a_{d},$
and
$g=b_{0}v^{e}+b_{1}v^{e-1}+\cdots+b_{e},$
with $a_{0},b_{0}\neq 0$. The resultant $\mathrm{Res}\,(f,g)$ is the determinant
$\begin{vmatrix}a_{0}&a_{1}&a_{2}&\cdots&a_{d}&0&\cdots&0\\ 0&a_{0}&a_{1}&\cdots&a_{d-1}&a_{d}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots&\ddots&\ddots&\vdots\\ 0&0&0&a_{0}&\cdots&\cdots&a_{d-1}&a_{d}\\ b_{0}&b_{1}&b_{2}&\cdots&b_{e}&0&\cdots&0\\ 0&b_{0}&b_{1}&\cdots&b_{e-1}&b_{e}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots&\ddots&\ddots&\vdots\\ 0&0&0&b_{0}&\cdots&\cdots&b_{e-1}&b_{e}\end{vmatrix}$
which has $e$ rows of $a_{i}$ and $d$ rows of $b_{j}$.
Alternatively, if $\lambda_{1},\cdots,\lambda_{d}$ and $\mu_{1},\cdots,\mu_{e}$ are respectively the roots of $f,g$ in the algebraic closure of $K$, counted with multiplicities, then
$\mathrm{Res}\,(f,g)=a_{0}^{e}b_{0}^{d}\prod_{\stackrel{{\scriptstyle 1\leq i\leq d}}{{1\leq j\leq e}}}(\lambda_{i}-\mu_{j}).$
For the polynomial $f$, the discriminant $\mathrm{Disc}\,(f)$ is defined as
$\mathrm{Disc}\,(f)=a_{0}^{2d-2}\prod_{i<j}(\lambda_{i}-\lambda_{j})^{2}.$
Proposition 6.4.
The following assertions are true:
  1. (i)
    $\mathrm{Disc}\,(f)=(-1)^{d(d-1)/2}a_{0}^{2d-1}\prod_{i\neq j}(\lambda_{i}-\lambda_{j})=\frac{(-1)^{d(d-1)/2}}{a_{0}}\mathrm{Res}\,\left(f,\frac{\mathrm{d}f}{\mathrm{d}v}\right).$
  2. (ii)
    $\mathrm{Res}\,(f,g)=(-1)^{de}\mathrm{Res}\,(g,f)$.
  3. (iii)
    If $d=0$, then $\mathrm{Res}\,(f,g)=a_{0}^{e}$.
  4. (iv)
    If $h$ is another polynomial in $K[v]$, then
    $\mathrm{Res}\,(fg,h)=\mathrm{Res}\,(f,h)\mathrm{Res}\,(g,h).$
  5. (v)
    If $a_{0}=1$, then for any $h\in K[v]$, we have
    $\mathrm{Res}\,(f,g)=\mathrm{Res}\,(f,g-fh).$
  6. (vi)
    If $f=f_{1}\cdots f_{k}$, then
    $\mathrm{Disc}\,(f)=\left(\prod_{i=1}^{k}\mathrm{Disc}\,(f)\right)\left(\prod_{1\leq i<j\leq k}\mathrm{Res}\,(f_{i},f_{j})^{2}\right).$
6.5. Intersection multiplicities
Now we use the Puiseux series and resultants to study intersection multiplicities of plane curve germs. Suppose $F,G$ are two reduced curve germs in $(\mathbb{C}^{2},0)$, with isolated singularity at 00, defined by
$F:\ f(u,v)=0,$
and
$G:\ g(u,v)=0.$
By the Weierstrass preparation theorem (see [6, Thm.1.8.7]), we may assume that $f,g$ are Weierstrass polynomials:
$f=v^{d}+a_{1}(u)v^{d-1}+\cdots+a_{d}(u),$
and
$g=v^{e}+b_{1}(u)v^{e-1}+\cdots+b_{e}(u),$
where $a_{i},b_{j}\in\mathbb{C}\{u\}$ satisfy $a_{i}(0)=b_{j}(0)=0$ for $i\leq d,j\leq e$. For convenience, we will also regard $a_{i},b_{j}$ as elements of $\mathbb{C}\langle\langle u\rangle\rangle$ by the inclusion $\mathbb{C}\{u\}\subset\mathbb{C}\langle\langle u\rangle\rangle$.
The intersection multiplicity of $F$ and $G$ at 00, denoted by $i_{0}(F,G)$, which is equal to the intersection multiplicity of $f$ and $g$ at 00, denoted by $i_{0}(f,g)$, is defined as
$i_{0}(F,G)=i_{0}(f,g)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{(f,g)}.$
The Milnor number of $F$ at 00, denoted by $\mu_{0}(F)$ or $\mu_{0}(f)$, is defined as the intersection multiplicity at 00 of $\frac{\partial f}{\partial u}$ and $\frac{\partial f}{\partial v}$:
$\mu_{0}(F)=i_{0}\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)}.$
The following results are classical; the interested reader may consult [6, Chap.2] or [9, Chap.I, §3].
Proposition 6.6.
  1. (i)
    $i_{0}(f,g)<\infty$ if and only if $F$ and $G$ do not share a common irreducible component.
  2. (ii)
    If $h$ is another element in $\mathbb{C}\{u,v\},h(0,0)=0$, then $i_{0}(fg,h)=i_{0}(f,h)+i_{0}(g,h)$.
  3. (iii)
    If $F$ is irreducible and $s=s(u)$ is a $v$-root of $f$ such that $\nu(s)=d$, then
    $i_{0}(f,g)=do_{u}\left(g(u,s(u))\right)=\sum_{\alpha=1}^{d}o_{u}\left(g\left(u,s(e^{2\pi\sqrt{-1}\alpha/d}u)\right)\right).$
  4. (iv)
    (Halphen’s formula) If $s_{1},\cdots,s_{d}$ and $s_{1}^{\prime},\cdots,s_{e}^{\prime}$ respectively are the $v$-roots of $f$ and $g$, then
    $i_{0}(f,g)=\sum_{\stackrel{{\scriptstyle 1\leq i\leq d}}{{1\leq j\leq e}}}o_{u}(s_{i}-s_{j}^{\prime}).$
    As a corollary, we have
    $i_{0}(f,g)=o_{u}\left(\mathrm{Res}\,(f,g)\right).$
  5. (v)
    If $f=f_{1}\cdots f_{k}$, then
    $\mu_{0}(f)=\sum_{i=1}^{k}\mu_{0}(f)+\sum_{i\neq j}i_{0}(f_{i},f_{j})-k+1.$
Now we establish the relationship between the Minor number and the discriminant.
Lemma 6.7.
With the notations above, we have
$o_{u}\left(\mathrm{Disc}\,(f)\right)=\mu_{0}(f)+d-1.$
Proof.
By [12, Lem.6.5.7, p.148], we have that the number
$i_{0}\left(pf+qu\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)$
is independent of $(p,q)$ provided $p,q\geq 0$ and $(p,q)\neq(0,0)$. It follows that
$o_{u}\left(\mathrm{Disc}\,(f)\right)=i_{0}\left(f,\frac{\partial f}{\partial v}\right)=i_{0}\left(u\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)=i_{0}\left(u,\frac{\partial f}{\partial v}\right)+\mu_{0}(f).$
Recall that $f=v^{d}+\sum a_{i}(u)v^{d-i}$. We get that
$i_{0}\left(u,\frac{\partial f}{\partial v}\right)=i_{0}(u,v^{d-1})=d-1.$
Hence, we get the desired equality $o_{u}\left(\mathrm{Disc}\,(f)\right)=\mu_{0}(f)+d-1$. ∎
6.8. Proof of Proposition 4.4
To give the proof of Proposition 4.4, we first recall our settings. We are given a reduced and irreducible curve $C$ on $X_{d}$, from which we get the following objects:
  1. (1)
    The branched Galois cover
    $\rho:\ X_{d}\to\mathbb{P}^{2},\quad[x,y,z,w]\mapsto[x,y,z].$
  2. (2)
    There exists a homogeneous polynomial $g(x,y,z)$ of degree $e$ such that
    $(g)=m(C_{1}+\cdots+C_{k}),$
    with $C_{i}^{2}=C^{2}$ for $i=1,\cdots,k$, and $k\geq 2$.
  3. (3)
    The subscheme $\mathscr{C}$ on $X_{d}$ defined by $g$, i.e.,
    $\mathscr{C}:\ g(x,y,z)=0.$
  4. (4)
    The curves $D$ and $E$ in $\mathbb{P}^{2}$ defined as
    $D:\ x^{d}+y^{d}+z^{d}=0\quad\mathrm{and}\quad E:\ g(x,y,z)=0.$
    In addition, $D$ and $E$ intersect at only finitely many points.
  5. (5)
    Fix an intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of different irreducible components of $\mathscr{C}$ satisfying $w_{0}=0$. We assume $x_{0}\neq 0$ and $z_{0}\neq 0$, and denote $p=[x_{0},y_{0},z_{0}]$.
  6. (6)
    Introduce the affine coordinates
    $u=\frac{y}{x}-\frac{y_{0}}{x_{0}},\qquad v=\frac{z}{x}-\frac{z_{0}}{x_{0}},\qquad\omega=\frac{w}{x},$
    and then the scheme $\mathscr{C}:g=0$ on $X_{d}$ is defined in $\mathbb{C}^{3}=\{x\neq 0\}\subset\mathbb{P}^{3}$ by
    $h(u,v)=0,\quad l(u,v)+\omega^{d}=0,$
    where
    $h(u,v)=g\left(1,u+\frac{y_{0}}{x_{0}},v+\frac{z_{0}}{x_{0}}\right)$
    and
    $l(u,v)=1+\left(u+\frac{y_{0}}{x_{0}}\right)^{d}+\left(v+\frac{z_{0}}{x_{0}}\right)^{d}.$
  7. (7)
    Viewed as an element in $\mathbb{C}\{u,v\}$, the function $h$ is $v$-general of order $b$.
Now we are ready to give the proof of Proposition of 4.4. To begin with, we first decompose $h$ in $\mathbb{C}\{u,v\}$ into irreducible factors
$h=h_{0}h_{1}\cdots h_{r};$
where $h_{1},\cdots,h_{r}$ are irreducible Weierstrass polynomial in $\mathbb{C}\{u\}[v]$ of order $b_{1},\cdots,b_{r}$ respectively, and $h_{0}$ is a nonzero constant. Recall that by assumption $h$ is $v$-general of order $b$, hence $b=b_{1}+\cdots+b_{r}$.
For $i\leq r$, let $s_{i}$ be a Puiseux series which is a root of $h_{i}$; without loss of generality, we assume $\nu(s_{i})=b_{i}$. Then, by Proposition 6.2 (iii) we have
$h_{i}=\prod_{\alpha=1}^{b_{i}}\left(v-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\ i\leq r.$
Set
$\phi_{i}(u,\omega)=\prod_{\alpha=1}^{b_{i}}\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right),\ i\leq r,$
and
$\Phi(u,\omega)=\phi_{1}\cdots\phi_{r}.$
Then $\Phi$ is a Weierstrass polynomial in $\omega$ of order $bd$ and moreover, the scheme $\mathscr{C}$ is locally defined by the equation $\Phi(u,\omega)=0$ in the local coordinate system $(u,\omega)$ on $X_{d}$.
(A) To show that $\mathscr{C}$ is reduced, it is sufficient to show
  1. ($a_{1}$)
    For $i\leq r$, $\phi_{i}$ is reduced; and
  2. ($a_{2}$)
    For $1\leq i<j\leq r$, $\phi_{i}$ and $\phi_{j}$ are coprime as elements in $\mathbb{C}\langle\langle u\rangle\rangle[\omega]$.
For ($a_{1}$), by definition of the intersection multiplicity, we have
$o_{u}\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)=\frac{i_{0}(h_{i},l)}{b_{i}}\leq\frac{i_{p}(D,E)}{b_{i}}<\infty.$
Therefore, $l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\neq 0$ and thus $\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)$ is reduced. Moreover, recall that
$l(u,v)=1+\left(u+\frac{y_{0}}{x_{0}}\right)^{d}+\left(v+\frac{z_{0}}{x_{0}}\right)^{d},$
hence, for $\alpha\neq\beta$,
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)$
$=$ $\left(s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\left(d\left(\frac{z_{0}}{x_{0}}\right)^{d-1}+r(u)\right)$
$\simeq$ $s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)$
where $\lim\limits_{u\to 0}r(u)=0$ and $\simeq$ means equality up to multiplication by an invertible element. Since $\nu(s_{i})=b_{i}$ and $\alpha\neq\beta$, we have
$s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\neq 0,$
and thus
$\gcd\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)=1.$
Consequently, $\phi_{i}$ is reduced.
For ($a_{2}$), it suffices to show that for any $i\neq j$ and for all $\alpha\leq b_{i},\beta\leq b_{j}$,
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\neq l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right).$
Similar to the discussions above, we have
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)-l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)\simeq s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u).$
Recall that $h_{i}$ and $h_{j}$ are coprime, they do not have common roots in $\mathbb{C}\langle\langle u\rangle\rangle$. Therefore, $s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)-s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\neq 0$.
Remark 6.9.
Proposition 4.4 (A) also follows from the fact that for all $q\in E\setminus(E\cap D)$, the inverse image $\rho^{-1}(q)\subset\mathscr{C}$ consists of exactly $d$ points.
(B) To compute $\mu_{P}(\mathscr{C})$, we use Lemma 6.7. Then
$\mu_{P}(\mathscr{C})=o_{u}\left(\mathrm{Disc}\,(\Phi)\right)-bd+1.$
Hence, we need to calculate $o_{u}\left(\mathrm{Disc}\,(\Phi)\right)$. By Proposition 6.4 (vi), we have from $\Phi=\phi_{1}\cdots\phi_{r}$ the following equality
$o_{u}\left(\mathrm{Disc}\,(\Phi)\right)=\sum_{i=1}^{r}o_{u}\left(\mathrm{Disc}\,(\phi_{i})\right)+\sum_{1\leq i<j\leq r}o_{u}\left(\mathrm{Res}\,(\phi_{i},\phi_{j})^{2}\right).$
Moreover, for $i\leq r$,
$o_{u}\left(\mathrm{Disc}\,(\phi_{i})\right)$
$=$ $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)$
$+\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right).$
We will prove the following results.
  1. ($b_{1}$)
    For $i\leq r$,
    $\sum_{\alpha=1}^{b_{i}}o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)=(d-1)i_{0}(h_{i},l),$
  2. ($b_{2}$)
    For $i\leq r$,
    $\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right)$
    $=$ $do_{u}\left(\mathrm{Disc}\,(h_{i})\right)$
    $=$ $d\left(\mu_{0}(h_{i})+b_{i}-1\right).$
  3. ($b_{3}$)
    For $1\leq i<j\leq r$,
    $o_{u}\left(\mathrm{Res}\,(\phi_{i},\phi_{j})^{2}\right)=do_{u}\left(\mathrm{Res}\,(h_{i},h_{j})^{2}\right).$
  4. ($b_{4}$)
    We have
    $o_{u}\left(\mathrm{Disc}\,(\Phi)\right)=(d-1)i_{p}(D,E)+d\left(\mu_{p}(E)+b-1\right).$
For ($b_{1}$), we have
$\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$ $\simeq$ $\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),w^{d-1}\right)$
$\simeq$ $\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{d-1};$
hence,
$o_{u}\left(\mathrm{Disc}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)\right)$ $=$ $(d-1)\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$
$=$ $\frac{(d-1)i_{0}(h_{i},l)}{b_{i}}.$
Summing over $\alpha=1,\cdots,b_{i}$ gives the desired equality.
For ($b_{2}$), we have
$\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)$
$=$ $\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)$
$\simeq$ $\left(l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{d}.$
Recall that
$l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\simeq s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u).$
Hence,
$\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)$
$\simeq$ $\left(s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{d}.$
It follows that
$\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)\right)\right)^{2}\right)$
$=$ $d\sum_{1\leq\alpha<\beta\leq b_{i}}o_{u}\left(s_{i}(e^{2\pi\sqrt{-1}\beta/b_{i}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{2}$
$=$ $do_{u}\left(\mathrm{Disc}\,(h_{i})\right)$
$=$ $do_{u}\left(\mu_{0}(h_{i})+b_{i}-1\right),$
where the last equality follows from Lemma 6.7.
For ($b_{3})$, we have
$\mathrm{Res}\,(\phi_{i},\phi_{j})^{2}$
$=$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),\omega^{d}+l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\mathrm{Res}\,\left(\omega^{d}+l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right),l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\right)^{2d}.$
Recall that
$l\left(u,s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)\right)-l\left(u,s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)\simeq s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u);$
hence,
$\mathrm{Res}\,(\phi_{i},\phi_{j})^{2}$
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(s_{j}(e^{2\pi\sqrt{-1}\beta/b_{j}}u)-s_{i}(e^{2\pi\sqrt{-1}\alpha/b_{i}}u)\right)^{2d}$
$\simeq$ $\mathrm{Res}\,(h_{i},h_{j})^{2d}.$
It follows that
$o_{u}\left(\mathrm{Res}\,(\phi_{i},\phi_{j})^{2}\right)=do_{u}\left(\mathrm{Res}\,(h_{i},h_{j})^{2}\right),$
as desired.
For ($b_{4}$), we combine ($b_{1}$)–($b_{3}$) and get
$o_{u}\left(\mathrm{Disc}\,(\Phi)\right)$ $=$ $\sum_{i=1}^{r}\left((d-1)i_{0}(h_{i},l)+do_{u}\left(\mathrm{Disc}\,(h_{i})\right)\right)+\sum_{1\leq i<j\leq r}\left(do_{u}\left(\mathrm{Res}\,(h_{i},h_{j})^{2}\right)\right)$
$=$ $(d-1)\sum_{i=1}^{r}i_{0}(h_{i},l)+do_{u}\left(\mathrm{Disc}\,(h)\right).$
Recall that $i_{0}(h,l)=i_{p}(E,D)$; and by Lemma 6.7, we have
$o_{u}\left(\mathrm{Disc}\,(h)\right)=\mu_{p}(E)+b-1.$
The equality in ($b_{4}$) follows immediately.
Now we are ready to finish the proof of (B). From ($b_{4}$) and the equality
$o_{u}\left(\mathrm{Disc}\,(\Phi)\right)=\mu_{P}(\mathscr{C})+bd-1,$
we get
$\mu_{P}(\mathscr{C})$ $=$ $(d-1)i_{p}(D,E)+d\left(\mu_{p}(E)+b-1\right)-bd+1$
$=$ $(d-1)i_{p}(D,E)+d\mu_{p}(E)-d+1.$
(C) Since $C_{1},\cdots,C_{k}$ are the irreducible components of $\mathscr{C}$, we have
$\mu_{P}(\mathscr{C})=\sum_{i\neq j}i_{P}(C_{i},C_{j})+\sum_{i=1}^{k}\mu_{P}(C_{i})-k+1.$
Note that $P$ is fixed under the Galois group for the branched cover $\rho$, and the Galois group acts transitively on $\{C_{1},\cdots,C_{k}\}$, hence
$\mu_{P}(C_{i})=\mu_{P}(C),\ \mathrm{for\ all\ }i=1,\cdots,k.$
It follows that
$\sum_{i\neq j}i_{P}(C_{i},C_{j})=\mu_{P}(\mathscr{C})-k\mu_{P}(C)+k-1.$
Using (B), we get
$\sum_{i\neq j}i_{P}(C_{i},C_{j})=(d-1)i_{p}(D,E)+d\mu_{p}(E)-k\mu_{P}(C)-(d-k).$
Summing over all possible intersection points and using Lemma 4.2, we obtain
$\sum_{i\neq j}C_{i}\cdot C_{j}=(d-1)D\cdot E+d\sum_{p\in D\cap E}\mu_{p}(E)-k\sum_{p\in D\cap E}\mu_{\rho^{-1}(p)}(C)-(d-k)\#(D\cap E).$
Note that $D\cdot E=de$, the equality in (C) follows immediately.
The proof of Proposition 4.4 is complete now.